Classwork: This worksheet may be used on as a reference during your exam. Production Curves ef Two Countries The coustries of Engl trade England, would prodace at land and haly have the production possibiities tables for sheep and hogs shown in the tables below, Without Albemative D and Italy woakd albo predace Alhermalive D. Note that the costs of producing sheep and bops are cosstant is both countries. Englands PRODUCTION POssBTes TABLE Assuse there are only two countries in the world. ys PRODUCTION POSSIBILITIES TABLE 16 128 G.5 Gs Task /. In England, the cost ofprodaang l sheep is: Tasł 2: In Italy, the cost of producing 1 sheep is: Task 3: Which country has the comparative advantage in sheep production and which country has the comparative advantage Ψ-2ols: 3 s,ME、abt tte hog prodaction? Englons Task: If each nation specializes in the product where it has a comparative advantage and trades with the otber, what will be the limits to the terms of trade for each sheepunt?./у 1-8) Tast 5 If the nations do not specialize and trade but remain at alternativse D in England and D in Italy, the combined production of Task &However, if the two nations specialize, the combined production of both countries will be how many sheep and how many Task 7: What will be the total gain of sheep and hogs if the countries specialize and trade? both countries will be how many sheep and how many hogs? bogs?

Need the Rest which I am lost in
Answer #1

England has a comparative advantage in hogs and Italy has a comparative advantage in sheeps.

5) At D, England produces 10 sheeps and 15 hogs, and Italy produces 8 sheeps and 9 hogs.

Combined production of sheeps= 10+8= 18

Combined production of hogs= 15+9= 24

6) with specialisation England produces 25 hogs only and Italy produces 20 sheeps only.

Combined production of sheeps= 20

Combined production of hogs= 25

7) total gains of sheeps= 20-18= 2

Total gains of hogs= 25-24= 1